### v2

# Self-Alignment: Improving Alignment of Cultural Values in LLMs via In-Context Learning

Abstract
Improving the alignment of Large Language Models (LLMs) with respect to the cultural values that they encode has become an increasingly important topic. In this work, we study whether we can exploit existing knowledge about cultural values at inference time to adjust model responses to cultural value probes. We present a simple and inexpensive method that uses a combination of in-context learning (ICL) and human survey data, and show that we can improve the alignment to cultural values across 5 models that include both English-centric and multilingual LLMs. Importantly, we show that our method could prove useful in test languages other than English and can improve alignment to the cultural values that correspond to a range of culturally diverse countries.
1 Introduction
The wide adoption of Large Language Models (LLMs) opened up pertinent questions as to how to correctly align LLM responses to reflect human intents and values (Cao et al., 2023; Shen et al., 2023; Liu et al., 2023; Xu et al., 2024; Wolf et al., 2023). While LLMs are already used by the public globally, much existing research shows that they are misaligned with respect to the cultural values that they encode (Arora et al., 2023; Choenni et al., 2024) and tend to exhibit western-centric biases (Naous et al., 2023). In practice though, to enable their adequate deployment in different languages, LLM output needs to be sensitive to the biases of the culturally diverse communities in which those languages are spoken (Hershcovich et al., 2022). Yet, popular alignment methods require large human preference datasets for fine-tuning and considerable computational resources (Rafailov et al., 2024; Ziegler et al., 2019), which makes it difficult and expensive to scale them to a multitude of languages and cultures. In this paper, we explore whether in-context learning (ICL), i.e., the LLMs’ ability to learn from a few demonstration examples at inference time (Wei et al., 2022; Brown et al., 2020; Dong et al., 2022), can be exploited to adjust the cultural values reflected in the LLM output, when provided with culturally-aligned demonstration examples. To the best of our knowledge, our work is the first to investigate this question.
The intuition behind ICL is that by providing more context to a given prompt, the model is able to pick up on cues in the demonstration examples and, consequently, adjust its responses accordingly without the need for gradient updates. In this paper, we test the ability of LLMs to adjust their responses to a given prompt based on demonstration examples exhibiting culture-specific values. Our hypothesis is that a set of statements that convey the cultural values of a particular country’s population would help to instantiate a particular cultural profile within the model, thus leading to a more culturally-aligned response. We refer to this technique as self-alignment.
Our demonstration examples are based on the questions from the World Values Survey (WVS) (Haerpfer et al., 2022), a social science effort documenting cultural values of participants in different countries. Specifically, we use the multilingual dataset of cultural value probes constructed by Arora et al. (2023) on the basis of the WVS. We use these probes both as prompts to probe the LLMs for their encoding of cultural values, and as our demonstration examples to evoke a cultural profile. Given that the probes are based on real survey questions, we can complete the demonstration examples with the real answer reported per country. See Figure 1 for an example.
The possible success of this method hinges on two main criteria: (1) it requires strong ICL capabilities to have emerged within the LLM, and (2) it requires the model to already encode associations between different cultural values such that it can detect cultural profiles and correct previously misaligned responses. Conducting experiments on a range of languages, we show that our self-alignment method improves the alignment of model responses to cultural value probes across 5 popular LLMs that include both English-centric and multilingual models. Moreover, we show that this success is not limited to English and the US values (the most commonly studied setting), but can also improve alignment of model responses in different languages and to the corresponding countries’ values, albeit to a different extent.
2 Related work
2.1 Misalignment of LLMs
While most popular LLMs have already undergone alignment in the form of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) at the fine-tuning stage, various studies show that LLMs are still not adequately aligned to a wide range of human values. For instance, Santurkar et al. (2023) compared model opinions with human responses in public opinion polls among various demographic groups and found substantial positional misalignment. Durmus et al. (2023) expanded this study to a global scale using cross-national surveys and found a bias towards Western countries, as well as unwanted cultural stereotypes. He et al. (2024) instead studied affective alignment, and measured how the emotional and moral tone of LLMs represents those of different groups. Finally, Arora et al. (2023); Choenni et al. (2024) and Cao et al. (2023) find that cultural values that LLMs encode in different languages do not align with human survey data, suggesting cultural misalignment.
2.2 Improving alignment at inference time
While there is a general consensus that LLMs should align to human values, such values vary considerably across countries, regions and even individuals. As such, it is underspecified what exactly we should aim to align the model to (Yao et al., 2023; Kirk et al., 2023). As a one-fits-all approach seems unlikely to lead to a satisfying outcome, and collecting a large alignment dataset for each scenario is prohibitively expensive, various approaches to improving alignment at inference time have been proposed. The benefit of such methods is that they would allow us to flexibly change alignment on an individual basis with minimal cost.
One such line of research focuses on sociodemographic prompting, a method to steer LLM responses towards answers that a persona, i.e., a human with a specific sociodemographic profile (e.g., age, gender, educational background, etc.), would give (Deshpande et al., 2023; Santurkar et al., 2023; Hwang et al., 2023; Cheng et al., 2023; He et al., 2024; Beck et al., 2024). While the methods used in these studies are also prompt-based, we exploit the ICL abilities of LLMs to trigger cultural profiles and instead improve alignment by inducing cultural knowledge into LLMs.
Kovač et al. (2023) are the first to study the perspective controlability in LLMs and introduce the notion of LLMs as superpositions of cultural perspectives. Their results show that prepending context to English prompts can induce different perspectives, leading to the question of how we can control such perspective changes. In this study, we delve deeper into this question and present ICL as a method for cultural perspective control. Moreover, we are the first to test cultural perspective controllability in a multilingual context. Finally, Sun et al. (2024) also use ICL to improve alignment, however, they do not explore cultural value alignment, and they use ICL to show a set of high quality responses that the model should mimic (as is traditionally done), but not to elicit a different perspective from the model.
3 Dataset
World Values Survey (WVS)
We aim to improve the alignment of cultural values for different language/country combinations using demonstration examples constructed from the World Values Survey (WVS) data (Haerpfer et al., 2022). More specifically we use the cloze-style probing templates, created by Arora et al. (2023), based on the WVS data. The WVS collects data on cultural values in different countries in waves, and our questions come from Wave 7 which ran from 2017 to 2020 and targeted 57 countries 111https://www.worldvaluessurvey.org. Survey results are published per question, organised in 13 categories: (1) Social Values, Attitudes and Stereotypes, (2) Happiness and Well-being, (3) Social Capital, Trust and Organisational Membership, (4) Economic Values, (5) Corruption, (6) Migration, (7) Security, (8) Post-materialist Index, (9) Science and Technology, (10) Religious Values, (11) Ethical Values and Norms, (12) Political Interest and Political Participation, (13) Political Culture and Regimes. Categories (4) and (8) are excluded as their questions could not be converted into probes. We use 237 probes which always prompt the model to choose between two answers (e.g. important and unimportant or agree and disagree) for templates such as: ’Religion is _ to me.’ and ’I _ that when a woman works for pay, the children suffer.’.
Multilingual probes
Arora et al. (2023) translated the English WVS probes into 13 languages: Romanian, Greek, Urdu, Farsi, Tagalog, Indonesian, German, Malay, Bengali, Serbian, Turkish, Vietnamese and Korean. These languages were selected using three criteria: (1) the languages can be mapped to one country covered by the WVS, (2) they are the official languages of the countries that they are mapped to, (3) the distribution of the language’s speakers can be primarily localized to a country or relatively small geographical region, and (4) all languages have at least 10K articles on Wikipedia such that the LLMs can have seen a sufficient amount of pretraining data. We use the multilingual probes to study how well our method performs in each language, when aligning to the cultural values of the country that this language is mapped to. For instance, we align LLMs in Romanian to the dominant responses for Romania from the WVS. See Appendix B for the full mapping.
4 Methods
4.1 Models
We test our method on 5 LLMs of varying sizes: Llama3-8B (Touvron et al., 2023), Mistral AI 7B (Jiang et al., 2023), CommandR 35B222https://docs.cohere.com/docs/command-r, Gemini-pro 1.5 50T333https://deepmind.google/technologies/gemini/ and BLOOMz 7B1 (Muennighoff et al., 2023). While most LLMs are English-centric, CommandR and BLOOMz were explicitly designed to be multilingual. For each LLM we use the instruction-tuned or chat fine-tuned version, see Appendix C for the full model details.
4.2 Prompt construction
We use the masked templates from Arora et al. (2023), but replace the [MASK] token with an underscore (see Figure 1 for an example of the prompt instructions). Moreover, we always complete each demonstration with the majority answer reported by the WVS results for the country that we are aligning to. Note that the WVS questions, in contrast to our probes, often ask for the degree to which the respondent agrees with the statement. As such, we aggregate the results from opposite ends of the scale to get a vote for each of our two classes (e.g. 1-5 is classified as disagree and 6-10 as agree). Finally, to ensure that the LLMs are not biased towards predicting one option over the other, we randomly pick which answer option to present first.
4.3 Demonstration selection strategies
We evaluate 4 different strategies for selecting 5 demonstration examples in the same language as the test example.
Fully random
We randomly select demonstration examples from the WVS dataset. While all examples are related to cultural values, the test and demonstration examples are not guaranteed to be of relevance to one another. Yet, as we always complete demonstrations with the cultural values from the same country, it could still provide a useful cue about the predominant cultural values of a country.
Random within category
We select random demonstration examples from the same WVS category as the test examples. The idea behind this is that because all examples come from the same category, the demonstrations should at least be thematically relevant to the test example.
ChrF++ scores within category
Besides selecting demonstration examples from the same WVS category, we use the chrF++ metric Popović (2017) to determine the similarity between each test example and all possible demonstrations. ChrF++ calculates the character and word -gram overlap between two strings. As such, we use it as an inexpensive method to select demonstrations that have a greater lexical overlap with the test example (e.g. ’Friends are important’, ’Family is important’ etc.).
ChrF++ scores across categories
We compute chrF++ scores between each test example and all demonstration examples in the WVS dataset. We do this to test the robustness of using chrF++ to find relevant demonstrations in the absence of categorical annotations in future cultural value datasets.
4.4 Evaluation
Due to the stochastic nature of LLMs during the sampling process, generated LLM responses to the same prompt can vary. A common pitfall in the evaluation of LLMs is that in order to try to force deterministic outputs, researchers tend to set the temperature value to zero. This is not a satisfying solution for two reasons: (1) It does not allow for a realistic evaluation of LLMs as in practice users will likely not change the hyperparameters to enforce determinism hence resulting in a mismatch between the model under study and the one used in practice, and (2) contrary to popular belief, setting the temperature to zero does not always guarantee deterministic output (Ouyang et al., 2023). Thus, we instead embrace the stochastic nature of LLMs and use the default hyperparameters (see Appendix C for details), but evaluate a distribution of model responses.
Comparing response distributions
For each prompt, we retrieve 10444Note that we also tested sampling 20 responses, but found that this did not change the results substantially. model responses in the zero-shot setting and when using self-alignment, which we refer to as the response distributions. For both response distributions, we then compute which percentage of answers is in line with the majority answer reported by the WVS survey. If the percentage of correct answers from the self-alignment distribution is higher, we consider the alignment improved. We then report for which percentage of test examples such an increase in alignment was detected when using self-alignment versus zero-shot prompts.
Reduction in error rate
Besides testing how many times our method was able to improve the alignment, we are also interested in analyzing how much the alignment tends to increase. Given that the initial misalignment varies per test example, we compute the reduction in error rate. For each test example for which we were able to improve the alignment, we compute how much the error decreased relative to the original misalignment:
where is the percentage of misalignment in the response distribution under the zero-shot setting and when using self-alignment.
4.5 Alignment procedure
Detecting misaligned examples
We prompt LLMs in each language without any demonstration examples in a zero-shot setting, and test how many answers from the response distribution correspond with the majority answer reported by the WVS. We then focus on the test examples for which the alignment is imperfect (<100 correct answers).
Value alignment
For each misaligned example, we attempt to adjust the alignment by prepending 5 demonstration examples in the same language as the test example, with their correct labels (according to the WVS result reported for the target country) to the original prompt, see Figure 1.
5 Self-alignment results in English
In Sections 5.1 and 5.2, we first test which demonstration selection strategy is most effective using Llama3-8B in English when aligning to the values reported by the WVS survey for the United States. In Section 5.3 we then test how the best strategy performs across models.
5.1 Differences in selection strategies
When testing Llama3-8B in English in the zero-shot setting, we identified 117/237 test examples that were misaligned555Due to randomness in response generations, the amount of misaligned examples can slightly differ across runs.. In Figure 2, we report the percentage of these misaligned test examples for which the alignment improves as a result of self-alignment. We find that chrF++ without restricting the selection to WVS categories gives the best performance. For 73 of test questions we can increase the alignment of the response distribution by prepending 5 demonstration examples to our prompt. As expected, we also find that random selection strategies underperform compared to using chrF++. This suggests that the content of the demonstration examples matters, and that the effectiveness of the method increases when the selected demonstration examples become more relevant to the test instance.
In Figure 3 we show to what extent the error rate tends to reduce when using our self-alignment method. From this, we observe that all strategies, apart from random, correct the alignment to a large extent. In fact, most of the corrections are centered around a 80-100 reduction in error. This indicates that our simple method is effective at fully correcting the response distribution.
5.2 Robustness analysis
Lu et al. (2022) have shown that the order in which demonstration examples are presented can impact the performance of LLMs. To test whether our method remains robust to different orderings of demonstration examples selected using chrF++, we, for each draw, randomly shuffle the examples before re-constructing the prompt. We find that the alignment still increases for 73 of the test examples. This suggests that the self-alignment method is not very sensitive to the order in which the demonstration examples are presented.
5.3 Generalisation across models
We found that self-alignment, coupled with chrF++ across categories as a demonstration selection strategy, can effectively improve cultural value alignment in Llama3-8B. We now test whether this result holds when using the same set up across a variety of LLMs. In Figure 4, we find that while performance is similarly high for Mistral and Gemini-pro, the method is somewhat less effective for CommandR and BLOOMz. However, while the percentage of misaligned examples for which alignment increases is lower for CommandR (65%), in most cases where it does not increase, the demonstration examples have no effect on the alignment at all (27%). We consider this a positive finding as one can argue that if alignment improves for some cases and (mostly) does not decrease for others, this still makes the method useful in practice.
Moreover, in Figure 6 we find that the reduction in error rate is again centered around 80-100. Yet, we find that for Llama3-8B, which increases the alignment for the most test examples (73), the reductions in error rates tend to be lower for more test examples (meaning that it improves the alignments less effectively). Interestingly, we find that CommandR and BLOOMz, where alignment increases for fewer test examples, reductions in error rates tend to 100 most often. Importantly, in Figure 5 we show that the distribution of the percentage of incorrect answers from the response distributions across test examples is relatively similar across models. As such, it is not much easier for BLOOMz and CommandR to achieve a higher reduction in error rate compared to the other LLMs 666Note that if the error rate is centered around e.g., 10, only one correct answer is needed to lead to a 100 reduction.. In fact, CommandR is the outlier as the percentage of misaligned answers from the response distribution is centered around 100, making it more difficult to fully correct. Taking both the amount of test examples for which alignment improves and the extent to which this happens into account, we find that self-alignment performs best on Mistral.
6 Self-alignment in diverse languages
We have seen that self-alignment can be effective. However, we have only tested the method in English and when aligning to US values. We now test how well this method works when using languages other than English. Note that for each model, we only test the languages that are supported by the LLM as we can not expect LLMs to understand how to align to cultures for which it did not see any pretraining data from the corresponding language.
| en | de | ro | ko | ur | id | vi | fa | el | tr | bn | sr | ms | tl | en | de | ro | ||
| CommandR | Llama3 | |||||||||||||||||
| 65 | 68 | 70 | 59 | - | 59 | 72 | 65 | 65 | 62 | - | - | - | - | 73 | 67 | 76 | ||
| 27 | 27 | 19 | 34 | - | 31 | 22 | 26 | 29 | 33 | - | - | - | - | 9 | 9 | 7 | ||
| 8 | 5 | 11 | 7 | - | 11 | 6 | 9 | 6 | 5 | - | - | - | - | 18 | 24 | 16 | ||
| Gemini-pro | Mistral | |||||||||||||||||
| 70 | 72 | 66 | 63 | - | 58 | 65 | - | 62 | 57 | 53 | 67 | - | - | 70 | 67 | 59 | ||
| 17 | 15 | 19 | 24 | - | 25 | 19 | - | 20 | 21 | 22 | 18 | - | - | 14 | 8 | 8 | ||
| 14 | 13 | 15 | 13 | - | 17 | 16 | - | 18 | 23 | 24 | 15 | - | - | 16 | 25 | 33 | ||
| BLOOMz | ||||||||||||||||||
| 51 | 60 | 63 | - | 49 | 51 | 59 | - | 69 | 68 | 53 | 55 | 56 | 65 | |||||
| 13 | 9 | 9 | - | 16 | 10 | 9 | - | 7 | 8 | 11 | 14 | 11 | 10 | |||||
| 31 | 31 | 28 | - | 36 | 39 | 32 | - | 24 | 24 | 36 | 32 | 32 | 25 | |||||
6.1 Initial alignment to cultural values
Before testing our method in different languages, we evaluate the initial alignment of the LLMs to the cultural values of the respective countries, in order to assess the degree of their misalignment. In Figure 7, we report the percentage of LLM responses that are misaligned across languages for each model/language combination. Note that we still apply strict criteria where examples that are not correctly answered across all 10 runs are classified as misaligned. We find that all models are relatively well aligned in English to US values compared to other language/country combinations. This is not surprising as LLMs are still predominantly trained on English data (Kew et al., 2023), and therefore tend to exhibit Western biases (Kotek et al., 2023; Adilazuarda et al., 2024). Moreover, we find that overall, BLOOMz exhibits the worst alignment to human values and this result holds across all test languages. This could explain the lower effectiveness of self-alignment reported for BLOOMz in percentages in Section 5.3, as there is a much larger number of misaligned examples in absolute numbers. Moreover, we find that BLOOMz also tends to be less consistent in its predictions, resulting in (almost) never being able to predict the right value 100 of the time. Moreover, across LLMs we find that test examples are especially poorly aligned in Romanian and Greek. For the distribution in error sizes across languages, see Appendix D. Overall, we find that these distributions are relatively similar across languages, except for CommandR.
6.2 Self-alignment results across languages
We now use the multilingual probes described in Section 3, and repeat the self-alignment experiments as before in different languages. The demonstration examples are selected using chrF++ from the dataset of the corresponding language. In Table 1, we find that self-alignment is effective across different languages. While the effectiveness on average drops slightly when using languages other than English, we also see many cases in which the method performs better on languages other than English. For instance, for Llama3-8B the method works best in Romanian (76), for CommandR in Vietnamese (72), for Gemini-pro in German (72), and for BLOOMz in Greek (69). Interestingly, we find that self-alignment is least biased to work well in English for BLOOMz (in 9/12 languages we are able to improve the alignment to a greater extent). This is interesting as it suggests that BLOOMz is not necessarily biased to adopt cultural values from the US more easily than for other countries. Importantly, BLOOMz is also the LLM for which the pretraining data was most balanced during training as it comprised of only 30 of English data. Moreover, in Section 6.1, we found that for BLOOMz the highest number of examples were misaligned, meaning that in absolute numbers, self-alignment on BLOOMz does not underperform compared to the other LLMs.
| Demonstrations | Q&A | Explanation | |||||||||||||||||
Further analysis
Given that the success of self-alignment appears to be dependent on the test language, we now study for which types of test examples we obtain alignment improvements across languages. To this end, for each language, we report the number of examples for which alignment improved in BLOOMz broken down per WVS category in Figure 8. We find that self-alignment tends to correct a similar numbers of examples for most WVS category across languages, but observe greater variation for the political categories. For instance, we find that test examples pertaining to political interest get particularly often corrected for Romanian and Vietnamese and to a much lesser extent for English. The low performance in English on this category could in general explain why higher scores were achieved by other languages. Moreover, we find other outliers, such as that for Turkish, relatively many examples pertaining to Religious values were corrected. Note that overall, we find similar trends across languages for all models, see Appendix E for results on the other LLMs.
7 Qualitative analysis
In previous sections, we quantitatively assessed the effectiveness of our method. We now perform a qualitative analysis to get a better idea of whether the LLM could be changing its predictions for the right reasons (Min et al., 2022). In particular, we repeat the experiments in English on misaligned examples, and this time, allow for more tokens to be generated such that the LLM can provide a local explanation to its generated answer post-hoc (Singh et al., 2024). In Table 2, we show three examples for which self-alignment led to improvements. In particular, we find that in most of such cases, the LLM explains its answer based on the patterns in the content of the demonstration examples. For instance, we find that the LLM distinguishes between confidence in different types of institutions (example 1), picks up on differences in attitudes towards positive and negative statements on immigration (example 2), and recognizes disagreement with gender-based generalizations and as such adjusts its responses to place men and women on the same footing (example 3).
Despite these promising findings, we also in some (rare) cases find that the model adjusts its responses based on spurious correlations (e.g., ‘Based on the previous examples, it seems that the pattern is to agree with the statement if previous statements were also agreed with, and disagree if previous statements were disagreed with.’). Future work should study (1) how faithful these generated explanations are (Chen et al., 2023; Ye and Durrett, 2022), and (2) how we can avoid the tendency of LLMs to exploit superficial patterns in the demonstration examples. In particular, this could reduce the cases for which we observe that alignment deteriorates instead of improves due to self-alignment.
8 Conclusion
In this paper, we presented self-alignment, a simple, yet novel, method to improve the alignment of LLMs at inference time with respect to the cultural values that they encode. Self-alignment exploits the ICL abilities of a LLM to adjust the model’s responses such that they better align to the cultural values of a country. We found that this method proves effective across 5 LLMs and across a variety of languages, albeit to a different extent. Moreover, we found evidence that the LLMs can indeed pick up useful cues from the demonstration examples to induce the correct answers. We envision that the ideas behind self-alignment could prove useful for LLMs in practice. In particular, the WVS survey value responses could be replaced by real user responses and automatically be prepended as demonstration examples to each prompt through the system message (Lee et al., 2024). However, future work should first study how this method will affect model responses in more realistic chat scenarios (i.e., when not explicitly prompting the LLM for cultural values).
9 Limitations
While this method shows promising results, this work is exploratory in that we were only able to evaluate the method in a somewhat artificial, controlled setting, where all samples were presented in the same format and were originally carefully curated by social scientists. In practice, however, users will not often explicitly ask the model to answer questions which directly probe for cultural values. Thus, while we show that ICL-based self-alignment is a promising approach in principle, future work should investigate to what extent these types of demonstration examples will cause the model to adjust its responses when talking about cultural values implicitly.
Moreover, as mentioned in the related work, what to align LLMs to is still an active topic of discussion. For this study, we always attempt to align the entire distribution of model responses to the majority answer from the WVS survey results. This is a sensible choice because we deliberately choose language and country pairs for which we can see the language as a proxy for culture (see Section 3). However, this might not necessarily be what we want in practice. It might also be sensible to aim for a soft alignment to better reflect the human data instead. For instance, we could allow the LLMs to be less certain about answers to questions for which human responses were more divided. When there is more consensus on what LLMs should be aligned to, this should also be taken into consideration.
10 Ethical considerations
All data sources used in this study are publicly available. While we test for cultural alignment to human data in this study, we recognize that languages can not simply be mapped to single countries and, therefore, it is not always straightforward to decide which human values the model should align to in practice. Human value alignment requires a nuanced approach, accounting for the values and cultural norms of diverse social groups (also within a country, society or linguistic realm). However, a full investigation of this is a major research direction and is out of scope of the current paper, which focuses on whether in-context learning can in principle provide a viable mechanism for LLM alignment at inference time, relying on existing datasets of cultural values from the social sciences.
Appendix A Language to country mapping
| Language | Country |
| English (en) | United States |
| Romanian (ro) | Romania |
| Greek (el) | Greece |
| Urdu (ur) | Pakistan |
| Farsi (fa) | Iran |
| Tagalog (tl) | Philippines |
| Indonesian (id) | Indonesia |
| German (de) | Germany |
| Malay (ms) | Malaysia |
| Bengali (bn) | Bangladesh |
| Serbian (sr) | Serbia |
| Turkish (tr) | Turkey |
| Vietnamese (vi) | Vietnam |
| Korean (ko) | South Korea |
Appendix B Model details
We use 5 LLMs of varying sizes of which all are open-source except for Gemini. For our open source model, we rely on the HuggingFace implementation (Wolf et al., 2019), and for Gemini we use Google’s paid API service.
Llama 3
Llama3-8B777https://ai.meta.com/blog/meta-llama-3/ is pretrained mostly on English data. While it covers data from some other languages, those languages are mostly limited to Indo-European languages written in the Latin script. We use the Meta-Llama-3-8B-instruct version with a temperature of 0.6 as proposed for this model in the HuggingFace documentation.
Mistral
Similar to Llama3-8B, Mistral AI 7B (Jiang et al., 2023) is English-centric and mostly pretrained on Latin-script languages. For our experiments, we use the Mistral-7B-Instruct-v0.2 checkpoint with the HuggingFace default temperature value of 1.0.
Gemini 1.5 Pro
Gemini 1.5 pro888https://deepmind.google/technologies/gemini/ 50T is a closed-source LLM. It is not fully clear which languages this Gemini version covers but it is reported to support over 35 languages. We query the gemini-pro version through Google’s official API with the default temperature value of 1.0. For running all our Gemini experiments, we spend 25 euros on API credits.
CommandR
We use Cohere’s CommandR 35B LLM999https://docs.cohere.com/docs/command-r that was optimized to perform well on English, French, Spanish, Italian, German, Brazilian Portuguese, Japanese, Korean, Simplified Chinese, and Arabic. In addition, the following 13 languages were seen during pretraining: Russian, Polish, Turkish, Vietnamese, Dutch, Czech, Indonesian, Ukrainian, Romanian, Greek, Hindi, Hebrew, Persian. We use the CohereForAI/c4ai-command-r-v01 checkpoint with a temperature of 0.3 as proposed for the model in the HuggingFace documentation.
BLOOMz
BLOOMz-7b1 was pretrained on 46 languages (Muennighoff et al., 2023). Importantly, BLOOMz is, unlike any other LLM, pretrained on many languages that are typically considered low-resource. In particular, many languages from the Indic and Niger-Congo family were included during pretraining, see https://huggingface.co/bigscience/bloom for the full list of pretraining languages. We use the bigscience/bloomz-7b1 checkpoint with the HuggingFace default temperature value of 1.0.
Appendix C Distribution of error sizes
Appendix D Alignment improvements across languages and WVS categories